it says, " Iddentify each curve, and find the characteristics listed."
(x+4)^2 + (y+7)^2 = 81 center,radius,eccentricity

Answer by lwsshak3(11628)

Iddentify each curve, and find the characteristics listed."
(x+4)^2 + (y+7)^2 = 81 center,radius,eccentricity
**
This is an equation of a circle.
Its standard form: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
For given equation:(x+4)^2 + (y+7)^2 = 81
center: (-4,-7)
radius=√81=9
eccentricity does not apply to circles